### v1

# Holographic dictionary for Lifshitz and hyperscaling violating black holes

Abstract
We develop a novel holographic dictionary for the thermodynamics of black holes with Lifshitz and hyperscaling violating asymptotics, generalizing the dictionary for Anti-de Sitter black holes. Using our dictionary we show that the holographic Euler equation is dual to a generalized Smarr formula for these black holes, and we find a precise match between the extended bulk and boundary first law. Notably, the dictionary for the central charge appearing in the Euler relation depends on the hyperscaling violating parameter, but not on the Lifshitz dynamical exponent.
preprint: APS/123-QED
Introduction. Holography, or gauge/gravity duality, is the idea that a theory of quantum gravity in a ($d+1$)-dimensional bulk spacetime is dual to a quantum theory without gravity living on the $d$-dimensional asymptotic boundary of the bulk spacetime [1, 2]. This duality is best understood for gravitational theories in asymptotically Anti-de Sitter (AdS) spacetime, for which the dual theory is a conformal field theory (CFT), also known as the ‘AdS/CFT correspondence’ [3, 4, 5]. An important feature of this correspondence is that the thermodynamics of stationary AdS black holes [6] is dual to the thermodynamics of high-energy equilibrium states in a strongly coupled, large-$N$ conformal gauge theory [7]. For instance, the black hole entropy $S=A/4G$ [8, 9] is identified with the thermodynamic entropy $\tilde{S}$ of the CFT equilibrium state, and the internal energy $\tilde{E}$, temperature $\tilde{T}$, electric potential $\tilde{\Phi}$ and charge $\tilde{Q}$ of the CFT state are proportional to, respectively, the mass $M$, temperature $T$, electric potential $\Phi$ and charge $Q$ of the black hole [10, 11, 12, 13]
$\tilde{E}=\frac{ML}{R}\,,\,\,\tilde{T}=\frac{TL}{R}\,,\,\,\tilde{S}=\frac{A}{4G}\,,\,\,\tilde{\Phi}=\frac{\Phi}{R}\,,\,\,\tilde{Q}=QL\,,$ (1)
where $R$ is the curvature radius of the spatial geometry on which the CFT lives, $L$ is the AdS curvature radius, $A$ is the black hole horizon area, and $G$ is Newton’s constant. The proportionality factors in the holographic dictionary for $\tilde{E},\tilde{T}$ and $\tilde{\Phi}$ appear due to a specific choice of CFT metric, for which the CFT time is $R/L$ times the static AdS time and the curvature radius is $R$ [4, 5, 14].
The general principles behind holography are expected to apply to non-asymptotically AdS spacetimes as well. Since the isometries of the bulk spacetime should match with the global symmetries of the dual boundary theory, field theories that are not conformally (or scale) invariant are not dual to AdS spacetime, but possibly to a different bulk geometry. For example, some quantum field theories exhibit the anisotropic scaling symmetry $\{t,x^{i}\}\to\{\alpha^{z}t,\alpha x^{i}\}$, parametrized by the dynamical critical exponent $z$. Together with time and space translations and space rotations, this scaling symmetry forms the Lifshitz symmetries. For $z\neq 1$, Lifshitz invariant theories describe non-relativistic quantum critical systems. The anisotropic scaling symmetry can be geometrically realized in the bulk by the so-called Lifshitz spacetimes, which are a generalization of AdS spacetime [15, 16, 17, 18] (see [19] for a review). Another generalization is given by hyperscaling violating geometries, which are dual to theories violating the hyperscaling property that entropy (or free energy) scales with its naive dimension. For hyperscaling violating Lifshitz (HVL) theories (on the plane), entropy scales with temperature as $S\propto T^{(d-1-\theta)/z}$, where $\theta$ is the hyperscaling violation exponent [20, 21]. Various black hole solutions with Lifshitz and/or hyperscaling violating asymptotics have been obtained in [18, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33].
In this note we develop a novel holographic dictionary for black hole thermodynamics in HVL theories. We find a generalization of the standard central charge dictionary $C\propto L^{d-1}/G$ for Einstein gravity, and obtain a dual relation between a generalized Smarr formula, relating the thermodynamic quantities in the bulk, and the holographic Euler relation on the boundary [11]
$\tilde{E}=\tilde{T}\tilde{S}+\tilde{\Phi}\tilde{Q}+\mu C\,,$ (2)
where the central charge $C$ is defined as the dimensionless proportionality factor of the grand canonical free energy, and $\mu$ is its conjugate chemical potential (we suppress the tildes on $\mu,C$ since they do not have an equivalent definition in the bulk). This Euler relation is valid for any large-$C$ field theory at finite temperature in the deconfined phase and it is the same for theories with different global symmetries (e.g., conformal and Lifshitz). Moreover, we show that an extended version of the first law for these black holes is dual to the thermodynamic first law on the boundary
$d\tilde{E}=\tilde{T}d\tilde{S}+\tilde{\Phi}d\tilde{Q}-pdV+\mu dC\,,$ (3)
provided that $\mu$ obeys the holographic Euler equation (2) and the boundary pressure $p$ satisfies the equation of state for Lifshitz and hyperscaling violating theories, see Eq. (28) below. Although we work within the context of a large class of static, charged analytic black hole solutions to a generalized Einstein-Maxwell-Dilaton (EMD) theory [33], we expect that our results apply to HVL black holes in general, thus showing that the recently developed holographic description of extended black hole thermodynamics in [11, 13] has a wide range of applicability.
Lifshitz and hyperscaling violating black holes. A line element of the general form [^1] Note that we are working with $d+1$ bulk spacetime dimensions, whereas [33] has $d+2$ bulk dimensions. Moreover, we redefined $V_{0}$ in [33] as $-2\Lambda_{0}$, and $\phi_{0}$ as $-\gamma\log r_{0}$, see (11).
$ds^{2}$ $=\chi\bigg{[}-\left(\frac{r}{L}\right)^{2z}f(r)dt^{2}+\frac{L^{2}}{f(r)r^{2}}dr^{2}+r^{2}d\Omega_{k,d-1}^{2}\bigg{]}\,,$
$\chi(r)$ $:=\left(\frac{r}{r_{F}}\right)^{-\frac{2\theta}{d-1}}\,,$ (4)
in $(d+1)$-dimensions, with $f(r)\rightarrow 1$ as $r\rightarrow\infty$, has been suggested to be dual to states in Lifshitz invariant field theories with hyperscaling violation. For $z=1$ and $\theta=0$ we recover asymptotically AdS spacetime. Here, $k$ parameterizes the topology of the $(t,r)=$ constant surfaces: $k=\{-1,0,1\}$ for hyperbolic, planar and spherical topology respectively. When $f(r)=1$ under the rescalings
$t\to\alpha^{z}t\,,\qquad\quad\Omega\to\alpha\Omega\,,\qquad\quad r\to\alpha^{-1}r\,,$ (5)
the line element (4) scales as
$ds\to\alpha^{\theta/(d-1)}ds\,,$ (6)
with $\alpha$ being a dimensionless scaling parameter. Leaving aside the scaling of the bulk radius $r$, this rescaling of the boundary metric is characteristic of hyperscaling violating field theories. Hence, we refer to (4) as asymptotically Lifshitz and hyperscaling violating geometries (although, strictly speaking, this is only true for $k=0$, since geometries with $k=\pm 1$ are not invariant under spatial translations.) As is standard for hyperscaling violating theories, the dual black hole geometry is considered to be valid up to a certain large radius $r_{F}$, which corresponds to a UV cut-off scale on the boundary, since the dual field theory could have a different UV completion [28, 31, 34].
[…]
$I$ $=\frac{1}{16\pi G}\int d^{d+1}x\sqrt{-g}\Big{[}R-\frac{1}{2}(\nabla\phi)^{2}+\mathcal{V}(\phi)$
$-\frac{1}{4}X(\phi)F^{2}-\frac{1}{4}Y(\phi)H^{2}-\frac{1}{4}Z(\phi)K^{2}\Big{]}\,,$ (7)
with field strengths $F=d\textsf{A},H=d\textsf{B}$ and $K=d\textsf{C}$ and
$\mathcal{V}=-2\Lambda_{0}e^{\lambda_{0}\phi}\,,\ X=X_{0}e^{\lambda_{1}\phi}\,,\ Y=Y_{0}e^{\lambda_{2}\phi}\,,\ Z=Z_{0}e^{\lambda_{3}\phi}\,,$
where the $\lambda_{i}$ are given in terms of ($d,\theta,z)$; namely:
$\lambda_{0}$ $=\frac{2\theta}{\gamma(d-1)}\,,\quad\lambda_{1}=-\frac{2(d-\theta-1+\theta/(d-1))}{\gamma}\,,$
$\lambda_{2}=$ $-\frac{2(d-2)(d-\theta-1)}{\gamma(d-1)}\,,\quad\lambda_{3}=\frac{\gamma}{d-\theta-1}\,,$
$\gamma=$ $\sqrt{2(d-\theta-1)\left(z-1-\frac{\theta}{d-1}\right)}\,,$ (8)
allowing for a non-zero hyperscaling violating parameter $\theta$. The gauge field A is responsible for the anisotropic scaling symmetry, B supports non-trivial (spherical or hyperbolic) horizon topology, and C allows for non-zero electric charge. The analytic solution for these black holes is given by (4) with [33]
$f(r)$ $=1+k\frac{(d-2)^{2}}{(d-\theta+z-3)^{2}}\frac{L^{2}}{r^{2}}-\frac{m}{r^{d-\theta+z-1}}$
$\quad+\frac{q^{2}}{r^{2(d-\theta+z-2)}}\,,$ (9)
C $=-\frac{\rho_{3}}{d-\theta+z-3}r_{0}^{\gamma\lambda_{3}}r^{-(d-\theta+z-3)}dt\,,$ (10)
$\phi$ $=\gamma\log(r/r_{0})\;,\quad$ (11)
generalizing previously discovered solutions for charged spherical black holes with Lifshitz asymptotics [27] (to non-zero $\theta$) and charged black branes with arbitrary Lifshitz and hyperscaling violation parameters [29] (to non-trivial topology). The respective mass and charge parameters are $m$ and $q$, and
$L^{2}$ $=-\frac{(d-\theta+z-2)(d-\theta+z-1)}{2\Lambda_{0}r_{F}^{2\theta/(d-1)}}r_{0}^{2\theta/(d-1)}\,,$ (12)
$\rho_{3}^{2}$ $=\frac{2q^{2}(d-\theta-1)(d-\theta+z-3)}{Z_{0}L^{2z}r_{0}^{\gamma\lambda_{3}}}r_{F}^{2\theta/(d-1)}\,,$ (13)
with $L$ the bulk curvature radius (analogous to the AdS curvature radius), and $r_{0}$ an arbitrary length scale. This solution is only valid for $d-\theta+z-3>0$ and $\theta<d-1$.
Extended bulk thermodynamics. The thermodynamic quantities of the hyperscaling violating Lifshitz black hole solution were computed in [33] (see also, e.g., [27, 29, 35, 36, 37]). The Hawking temperature $T$, Bekenstein-Hawking entropy $S$, ADM mass $M$, electric charge $Q$, and electric potential $\Phi$ are [^2] The ADM mass was obtained in [33] through the background subtraction method, so that the solution with $m=0$ has zero total gravitational energy. For hyperbolic black holes, however, this solution does not correspond to the ground state of the grand canonical ensemble, which is given by the extremal solution with zero charge and in our conventions hence has negative energy (whereas in [33] the ADM mass for $k=-1$ was redefined such that the extremal solution has zero energy).
$T$ $=\frac{f^{\prime}(r_{h})}{4\pi}\Bigl{(}\frac{r_{h}}{L}\Bigr{)}^{z+1}\!\!=\frac{r_{h}^{z}}{4\pi L^{z+1}}\left[d-\theta+z-1+\frac{k(d-2)^{2}}{d-\theta+z-3}\frac{L^{2}}{r_{h}^{2}}-\frac{(d-\theta+z-3)q^{2}}{r_{h}^{2(d-\theta+z-2)}}\right]\,,\quad S=\frac{\omega_{k,d-1}}{4G}r_{h}^{d-\theta-1}r_{F}^{\theta}\,,$
$M$ $=\frac{\omega_{k,d-1}}{16\pi G}(d-\theta-1)mL^{-z-1}r_{F}^{\theta}\,,\quad Q=\frac{\omega_{k,d-1}}{16\pi G}Z_{0}\rho_{3}L^{z-1}r_{F}^{\theta-2\theta/(d-1)}\,,\quad\Phi=\frac{qr_{F}^{\theta/(d-1)}r_{0}^{\lambda_{3}\gamma/2}}{\sqrt{\frac{Z_{0}(d-\theta+z-3)}{2(d-\theta-1)}}L^{z}r_{h}^{d-\theta+z-3}}\,,$ (14)
where $r_{h}$ is horizon radius, given by the largest positive root of $f(r)=0$. The mass $M$ depends on the parameters $(m,L,r_{F})$; $m$ can be expressed, in turn, in terms of $r_{h}$, $L$, and $q$ by solving $f(r_{h},L,m,q)=0$ for $m$. To derive a generalized Smarr relation and first law, we express $M$ as a function of the thermodynamic variables $S$ and $Q$, and the analog of the “bulk pressure” $P:=-\frac{\Lambda_{0}}{8\pi G}$ for EMD theory [38], where $\Lambda_{0}$ is the “bare” cosmological constant appearing in the Lagrangian. Due to the coupling of $\Lambda_{0}$ to the dilaton, this does not correspond to the pressure of a bulk perfect fluid as usual for the cosmological constant, but we will continue to use $P$ as a useful theory parameter. The result (see Supplemental Material) is that $M=M(S,Q,P)$. The mass also depends on the length scales $r_{F}$ and $r_{0}$, but we treat them here as fixed parameters, since they do not have a clear thermodynamic interpretation from the bulk and boundary perspective.
From the mass function (35) the following scaling relation may be inferred:
$M(S,Q,P)=\alpha^{2+\theta-d}M(\alpha^{d-\theta-1}S,\alpha^{d-\theta+z-3}Q,\alpha^{-2}P)\,,$ (15)
which arises from the rescaling $r_{h}\to\alpha r_{h}$, $L\to\alpha L$ and $q\to\alpha^{d-\theta+z-2}q$. A generalized Smarr formula can now be derived from this scaling relation by applying Euler’s theorem for homogeneous functions [38]
$(d\!-\!\theta\!-\!2)M\!=\!(d\!-\!\theta\!-\!1)TS\!+\!2\Theta P\!+\!(d\!-\!\theta\!+\!z\!-\!3)\Phi Q\,,$ (16)
with
$\Theta$ $:=-\frac{\partial M}{\partial P}\Big{|}_{S,Q}$ (17)
the conjugate quantity to $P$. If $\theta=0$ and $z=1$, $\Theta$ reduces to the expected expression $-r^{d}\omega_{k,d-1}/d$. Hence $\Theta$ (see equation (42)) is the HVL generalization of (minus) the “thermodynamic volume”, see e.g. [39].
The generalized Smarr relation (16) is new in the literature, as the scaling relation (15) was not properly derived before. In fact, in [35, 33, 40] it was assumed that the standard Smarr formula for AdS black holes (given by (16) with $z=1$ and $\theta=0$) also holds HVL black holes. However this is clearly not the case, since the Smarr relation (16) explicitly involves $z$ and $\theta$.
Next, from the function $M=M(S,Q,P)$ a variational relation follows straightforwardly. Viewing $r_{F}$ and $r_{0}$ as fixed, the extended first law for charged HVL black holes reads
$dM=TdS+\Phi dQ-\Theta dP\,.$ (18)
Since $P$ depends via $\Lambda_{0}$ on $L$ due to (12), using (16) we can write the extended first law as
$dM=TdS+\Phi dQ+\!\Big{(}(d-\theta-2)M\!-\!(d-\theta-1)TS$
$\!-\!(d-\theta+z-3)\Phi Q\Big{)}\frac{dL}{L}$ (19)
This expression is useful for finding a precise match between the bulk and boundary first laws.
Holographic dictionary. We would now like to find a gauge/gravity dictionary matching (19) to the thermodynamic first law of the dual field theory, given by
$d\tilde{E}=\tilde{T}d\tilde{S}+\tilde{\Phi}d\tilde{Q}-pdV+\mu dC$ (20)
where $\tilde{E}$ is the internal energy of the field theory, $\tilde{S}=S$ and $\tilde{T}$ its entropy and temperature, $\tilde{\Phi}$ and $\tilde{Q}$ its electric potential and charge, $p$ and $V$ its pressure and volume, and $\mu$ and $C$ its chemical potential and central charge [11]. We assume that the gauge theory lives on a spatial geometry with line element $R^{2}d\Omega_{k,d-1}^{2}$, where $R$ is the boundary curvature scale, so that the spatial volume is proportional to
$V\propto R^{d-1}\,.$ (21)
The dictionary for the central charge can be read off from the proportionality constant of the grand canonical free energy $F=M-TS-\Phi Q$. Since the mass, entropy and charge all scale with the central charge, $M,S,Q\propto C$, we can directly determine its dictionary entry by looking at, for instance, the Bekenstein-Hawking entropy (Holographic dictionary for Lifshitz and hyperscaling violating black holes). The central charge is dimensionless and should only depend on theory parameters, not on the specific details of the state, such as its temperature. Therefore, the dictionary for $C$ cannot involve the horizon radius $r_{h}$, which is related to the temperature (Holographic dictionary for Lifshitz and hyperscaling violating black holes). Multiplying and dividing the entropy by $L^{d-\theta-1}$, we can write
$S\propto Cx^{d-\theta-1}\,,$ (22)
with $x=r_{h}/L$, where we identify the central charge as
$C\propto\frac{L^{d-\theta-1}r_{F}^{\theta}}{G}\,.$ (23)
For $\theta=0$ this dictionary reduces to the well-known result for Einstein gravity. It is interesting to note that the dictionary entry (23) depends on $\theta$ but not $z$. It is well known that $\theta$ effectively lowers the number of dimensions, which is also apparent from (23). Moreover, it is quite natural that (23) involves $r_{F}$, since that corresponds to a UV cut-off in the boundary theory, on which the number of field degrees of freedom depends. Likewise, the entropy does not depend on $r_{0}$ and so the entry (23) for $C$ does not, since $S\propto C$.
The dictionary for the internal energy and temperature follows from symmetry arguments. The products $ML$ and $TL$ in the bulk are dimensionless, and so should match with products of boundary quantities that are (from (5)) invariant under the anisotropic scaling symmetry, $\{t,x^{i}\}\to\{\alpha^{z}t,\alpha x^{i}\}$, and the simultaneous hyperscaling transformation $\{t,x^{i}\}\to\alpha^{-\theta/(d-1)}\{t,x^{i}\}$ that undoes the rescaling (6) of the metric at the boundary. The invariant combinations are
$\tilde{E}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}=ML\,,\quad\tilde{T}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}=TL\,.$ (24)
Note that the scaling dimension of energy and temperature is $[\tilde{E}]=[\tilde{T}]=-[t]=z-\theta/(d-1)$, and the scaling dimension of the boundary curvature scale and spatial coordinate is $[R]=[x^{i}]=-1+\theta/(d-1)$ (see Appendix D in [36]), so that the LHS of (24) is indeed dimensionless.
The dictionary entries for the electric potential and charge are more complicated, but also essentially follow from dimensional analysis. We take the boundary electric charge $\tilde{Q}$ to be dimensionless, as for the entropy and central charge in (20); the boundary potential $\tilde{\Phi}$ then has the same dimensions as $\tilde{E}$ and $\tilde{T}$, so that $\tilde{\Phi}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}$ is scale invariant. The bulk charge scales as $Q\propto qL^{-1}r_{F}^{\theta-\theta/(d-1)}r_{0}^{1-z+\theta/(d-1)}/G$ and the bulk potential as $\Phi\propto qr_{F}^{\theta/(d-1)}r_{0}^{z-1-\theta/(d-1)}/(L^{z}r_{h}^{d-\theta+z-3})$, where the charge parameter and Newton’s constant have mass dimension $[q]=-d+\theta-z+2$ and $[G]=-d+1$, respectively. We can thus make the boundary charge $\tilde{Q}$ and the product $\tilde{\Phi}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}$ dimensionless with the following dictionary
$\tilde{\Phi}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}$ $=L^{z-1}r_{F}^{-\frac{\theta}{d-1}}r_{0}^{1-z+\frac{\theta}{d-1}}\Phi\,,$ (25)
$\tilde{Q}$ $=L^{2-z}r_{F}^{\frac{\theta}{d-1}}r_{0}^{z-1-\frac{\theta}{d-1}}Q\,,$ (26)
which satisfies
$\tilde{\Phi}\tilde{Q}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}=\Phi QL\,.$ (27)
Note this product is independent of $r_{0}$ and $r_{F}$. For $z=1$ and $\theta=0$ this agrees with the standard holographic dictionary [11]. There are no $\sqrt{G}$ factors due to a different normalization of the action compared to that in [13] (in our conventions the gauge coupling constant depends on $G$). We further comment that this corrects an earlier dictionary (Appendix D in [11]) for the electric charge and potential of Lifshitz black holes, where it was proposed that $\tilde{\Phi}=\Phi/R^{z}$ and $\tilde{Q}=QL$, which is different from the dictionary above for $\theta=0$, i.e. $\tilde{\Phi}=\Phi L^{z}/R^{z}$ and $\tilde{Q}=QL^{2-z}$. This difference ultimately arises from the fact that in [11] (following [35, 33]) it was assumed that the AdS Smarr formula also holds for Lifshitz black holes, whereas here we find that a more careful study of the scaling relation for the mass shows that the correct Smarr formula is (16).
Crucially, using the dictionary in (21) and (23)-(26) one can show that the extended bulk first law (19) exactly matches with the boundary first law (20). [^3] We also checked that the matching between the bulk and boundary extended first laws continues to hold if $r_{F}$ and $r_{0}$ are allowed to vary in the bulk (see Supplemental Material). The matching only works if the pressure satisfies the equation of state for HVL theories [19, 36]
$p=\frac{\tilde{E}}{V}\frac{z-\theta/(d-1)}{d-1-\theta}\,,$ (28)
and if the chemical potential $\mu$ obeys the holographic Euler equation (2). The equation of state (28) is a consequence of the scaling relation
$\tilde{E}(\tilde{S},\alpha^{d-1-\theta}V,\tilde{Q},C)=\alpha^{-z+\frac{\theta}{d-1}}\tilde{E}(\tilde{S},V,\tilde{Q},C)\,,$ (29)
which follows from the anisotropic scaling symmetry and the hyperscaling violating property. Differentiating (29) with respect to $\alpha$ and setting $\alpha=1$ yields (28), where the pressure is defined as $p:=-(\partial\tilde{E}/\partial V)_{\tilde{S},\tilde{Q},C}$.
On a planar boundary geometry (for $k=0$) there is an additional relation between the thermodynamic quantities, which is dual to an extra Smarr formula in the bulk. This allows us to replace the $\mu C$ term in the Euler relation with a $pV$ term, but this comes at the expense of additional $(z,\theta,d)$-dependent prefactors in front of the $pV$ and $\tilde{\Phi}\tilde{Q}$ terms (that are absent for $z=1,\theta=0$), namely:
$\tilde{E}=\tilde{T}\tilde{S}+\frac{d-\theta+z-2}{d-\theta-1}\tilde{\Phi}\tilde{Q}-\frac{z}{z-\theta/(d-1)}pV\,.$ (30)
We refer to the Supplemental Material for more details.
Finally, we demonstrate that the large-$C$ Euler equation (2) is dual to the generalized Smarr relation (16). First, we express the $\Theta P$ term in the Smarr formula in terms of the boundary energy $\tilde{E}$
$2\Theta P$ $=-2P\frac{\partial M}{\partial P}\Big{|}_{S,Q}=L\frac{\partial M}{\partial L}\Big{|}_{S,Q}$
$=R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}\frac{\partial\tilde{E}}{\partial L}\Big{|}_{S,Q}-\frac{\tilde{E}}{L}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}\,.$ (31)
In the first equality we used the definition of $\Theta$ (17), and in the second equality we employed the fact that $\Lambda_{0}$ depends on $L,r_{F}$ and $r_{0}$ via (12), but the latter two length scales are fixed. In the final equality we inserted the dictionary that relates $M$ and $\tilde{E}$. Subsequently, the partial derivative can be computed by noting that the boundary energy depends on the bulk quantities as follows:
$\tilde{E}=\tilde{E}(S,\tilde{Q}(L,Q),V(R),C(L))\,,$ (32)
yielding
$\frac{\partial\tilde{E}}{\partial L}\Big{|}_{S,Q}=\frac{1}{L}\left((2-z)\tilde{\Phi}\tilde{Q}+(d-\theta-1)\mu C\right)\,.$ (33)
To obtain this we used the dictionary for $\tilde{Q}$ (26) and $C$ (23). Inserting this back into (Holographic dictionary for Lifshitz and hyperscaling violating black holes) gives
$2\Theta P$ $=R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}\frac{1}{L}\left((2-z)\tilde{\Phi}\tilde{Q}+(d-\theta-1)\mu C-\tilde{E}\right)\,.$ (34)
This precisely agrees with the generalized Smarr formula (16) if we insert the large-$C$ Euler equation (2) and employ the holographic dictionary (24), (25) and (26). This extends previous derivations [10, 11, 13] of the Smarr formula from the Euler equation to HVL black holes, and it corrects Appendix D in [11] (see [41] for an alternative holographic interpretation of the Smarr formula). In the Supplemental Material we give further evidence for the large-$C$ Euler equation by deriving it from a generalized Cardy formula for $d=2$, without invoking insights from holography.
Conclusion. Using the extended black hole thermodynamics as a guiding principle, we developed a novel holographic dictionary for thermal equilibrium states dual to HVL black holes, and checked that the first law and large-$C$ Euler equation are still valid for these states. It is remarkable that the Euler equation has such a wide range of applicability.
For future work, it would be interesting to derive the dictionary for the thermodynamic black hole quantities using holographic renormalization [42, 43, 44] and, in particular, check our formula for the central charge (which can be read off from the Casimir energy for even $d$ and $k\neq 0$). Moreover, we would like to understand whether there exists a thermodynamic interpretation of the length parameters $r_{F}$ and $r_{0}$. Finally, it is likely that our new setup will lead to a discovery of so far unseen phase transitions and critical phenomena arising for Lifshitz and hyperscaling violating theories.
Acknowledgments. D.K. is grateful for support from GAČR 23-07457S grant of the Czech Science Foundation and the Charles University Research Center Grant No. UNCE24/SCI/016. M.R.V. acknowledges support from the SNF Postdoc Mobility grant P500PT-206877 and the Spinoza Grant of the Dutch Science Organisation (NWO). R.B.M. acknowledges support from the Natural Sciences and Engineering Research Council of Canada.
Supplemental Material
Appendix A Mass formula, scaling relations and extended first law
The formula for the ADM mass as a function $M=M(S,Q,P,r_{F},r_{0}^{-\gamma})$ is given by
$M$ $=(d-1-\theta)\frac{r_{F}^{\theta}\omega_{k,d-1}}{16\pi G}\bigg{(}\frac{16\pi GP}{(d-2+z-\theta)(d-1+z-\theta)}\bigg{)}^{\frac{z+1}{2}}\bigg{(}\frac{4GS}{\omega_{k,d-1}r_{F}^{\theta}}\bigg{)}^{\frac{d-1+z-\theta}{d-1-\theta}}\bigg{(}\frac{r_{0}}{r_{F}}\bigg{)}^{-\frac{(z+1)\theta}{d-1}}\times$
×$\quad\times\bigg{(}1+\frac{(d-2)^{2}(d-\theta+z-2)(d-\theta+z-1)k}{16\pi GP(d-\theta+z-3)^{2}}\left(\frac{r_{0}}{r_{F}}\right)^{\frac{2\theta}{d-1}}\left(\frac{4GS}{\omega_{k,d-1}r_{F}^{\theta}}\right)^{-\frac{2}{d-\theta-1}}$
$\qquad+\frac{8\pi GQ^{2}r_{0}^{2(z-1)}(d-\theta+z-2)(d-\theta+z-1)}{PZ_{0}r_{F}^{2\theta}(d-\theta-1)(d-\theta+z-3)\omega_{k,d-1}^{2}}\left(\frac{4GS}{\omega_{k,d-1}r_{F}^{\theta}}\right)^{-\frac{2(d-\theta+z-2)}{d-\theta-1}}\bigg{)}\,.$ (35)
In this appendix we treat the length scales $r_{0}$ and $r_{F}$ as variable parameters, and we derive scaling relations for the mass function that follow from rescaling $r_{0}$ and $r_{F}$. In the main text we kept $r_{0}$ and $r_{F}$ fixed, since they have no clear thermodynamic interpretation, but it is tantalizing to think that they might have a thermodynamic meaning in an ‘extended’ version of black hole mechanics (perhaps by varying the charges associated to the A and B gauge fields, see (3.13) and (3.14) in [33]).
Using (35) it is straightforward to directly verify the following three independent scaling relations:
$M(S,Q,P,r_{F},r_{0}^{-\gamma})\,$ $=\alpha^{-\theta}M(\alpha^{\theta}S,\alpha^{\theta-\theta/(d-1)}Q,\alpha^{-2\theta/(d-1)}P,\alpha r_{F},r_{0}^{-\gamma})\quad$ (38)
$=M(S,\alpha^{\lambda_{3}/2}Q,\alpha^{-\lambda_{0}}P,r_{F},\alpha r_{0}^{-\gamma})$
$=\alpha^{2+\theta-d}M(\alpha^{d-\theta-1}S,\alpha^{d-\theta+z-3}Q,\alpha^{-2}P,r_{F},r_{0}^{-\gamma})\,,$
which respectively arise from the rescalings ($r_{F}\to\alpha r_{F}$, $r_{0}^{-\gamma}\to\alpha r_{0}^{-\gamma}$), and ($r_{h}\to\alpha r_{h}$, $L\to\alpha L$ and $q\to\alpha^{d-\theta+z-2}q$). In AdS space ($z=1$, $\theta=0$) only the third scaling relation survives, from which the Smarr formula for AdS black holes is usually derived by applying Euler’s theorem for homogeneous functions [38]. For arbitrary $z$ and $\theta,$ however, we have two additional scaling relations due to the presence of two new length scales $r_{F}$ and $r_{0}.$ By applying Euler’s theorem to the three scaling relations (38)-(38), we thus respectively find three independent Smarr-type formulas for HVL black holes
$\theta M=\theta TS+\frac{(d-2)\theta}{d-1}\Phi Q+Ur_{F}+\frac{2\theta}{d-1}\Theta P\,,$ (39)
$0=\frac{\lambda_{3}}{2}\Phi Q+U_{2}r_{0}^{-\gamma}+\lambda_{0}\Theta P\,,$ (40)
$(d\!-\!\theta\!-\!2)M\!=\!(d\!-\!\theta\!-\!1)TS\!+\!2\Theta P\!+\!(d\!-\!\theta\!+\!z\!-\!3)\Phi Q\,.$ (41)
The conjugate variables to the “bulk pressure” $P:=-\frac{\Lambda_{0}}{8\pi G}$ and to the length scales $r_{F}$ and $r_{0}^{-\gamma}$ are defined as partial derivatives of the mass, and are, respectively, given by
$\Theta:=-\frac{\partial M}{\partial P}\Big{|}_{S,Q,r_{F},r_{0}}$ $=-\frac{L^{1-z}r_{F}^{\frac{(d+1)\theta}{d-1}}r_{0}^{-\gamma\lambda_{0}}r_{h}^{d-\theta+z-3}(d-\theta-1)\omega_{k,d-1}}{2(d-\theta+z-3)^{2}(d-\theta+z-2)(d-\theta+z-1)}\times$ (42)
×$\quad\times\bigg{(}r_{h}^{-2(d+z-1)}(d-\theta+z-3)^{2}\big{(}(z+1)r_{h}^{2(d+z)}+q^{2}(z-1)r_{h}^{2\theta+4}\big{)}+(d-2)^{2}kL^{2}(z-1)\bigg{)}\,,$
$U:=\frac{\partial M}{\partial r_{F}}\Big{|}_{S,Q,P,r_{0}}$ $=-\frac{\theta\omega_{k,d-1}L^{-z-1}r_{F}^{\theta-1}r_{h}^{-d-\theta-z-3}}{16\pi G(d-1)(d-\theta+z-3)^{2}}\bigg{[}r_{h}^{2}(d-\theta+z-3)^{2}\bigg{(}q^{2}r_{h}^{2\theta+4}((1-2z)d+\theta+(\theta+2)z-1)$
$\quad+r_{h}^{2(d+z)}((1+z)\theta+1-d)\bigg{)}-(d-2)^{2}kL^{2}r_{h}^{2(d+z)}(d+(1-z)\theta-1)\bigg{]}\,,$ (43)
$U_{2}:=\frac{\partial M}{\partial(r_{0}^{-\gamma})}\Big{|}_{S,Q,P,r_{F}}$ $=\frac{(d-\theta-1)\omega_{k,d-1}L^{-z-1}r_{0}^{\gamma}r_{F}^{\theta}r_{h}^{-d-\theta-z-3}}{32\pi G(d-\theta+z-3)^{2}}\bigg{[}(d-2)^{2}\lambda_{0}kL^{2}(z-1)r_{h}^{2(d+z)}$
$\quad+r_{h}^{2}(d-\theta+z-3)^{2}\left(\lambda_{0}(z+1)r_{h}^{2(d+z)}+q^{2}r_{h}^{2\theta+4}(\lambda_{0}(z-1)-2\lambda_{3})\right)\bigg{]}\,.$ (44)
The quantity $U$ vanishes if $\theta=0$, and $U_{2}$ is zero if both $\theta=0$ and $z=1$, in which case $\Theta$ reduces to $-r^{d}\omega_{k,d-1}/d$.
It is possible to get rid of the explicit $z$ and $\theta$ dependence in the Smarr relation, but this comes at the expense of introducing two additional terms, proportional to $r_{F}$ and $r_{0}^{-\gamma}$, yielding the generalized Smarr formula
$(d-2)M=(d-1)TS+(d-2)\Phi Q+2\Theta P+Ur_{F}-\gamma U_{2}r_{0}^{-\gamma}\,,$ (45)
obtained by combining (39)-(41). Note that the prefactors are the same as for the Smarr formula for AdS black holes [38]. Indeed, for $(\theta=0$, $z=1)$, $U=U_{2}=0$, reproducing the AdS Smarr formula. The Smarr relation (45) also follows from the scaling relation
$M(\alpha^{d-1}S,\alpha^{d-2}Q,\alpha^{-2(1+\theta/(d-1))}P,\alpha r_{F},\alpha^{-\gamma}r_{0}^{-\gamma})=\alpha^{d-2}M(S,Q,P,r_{F},r_{0}^{-\gamma})\,,$ (46)
which is a combination of the scaling relations (38)-(38).
Next, we use the function $M=M(S,Q,P,r_{F},r_{0}^{-\gamma})$ to derive an extended version of the first law for charged HVL black holes, that includes variations of $r_{F}$ and $r_{0}^{-\gamma},$
$dM=TdS+\Phi dQ-\Theta dP+Udr_{F}+U_{2}dr_{0}^{-\gamma}\,.$ (47)
Since the pressure depends via $\Lambda_{0}$ on $L$, $r_{F}$ and $r_{0}$, due to (12), the extended first law can also be written in terms of variations of these length scales
$dM=TdS+\Phi dQ+\big{(}\theta M-\theta TS-\frac{(d-2)\theta}{d-1}\Phi Q\big{)}\frac{dr_{F}}{r_{F}}$
$+\!\big{(}(d-\theta-2)M\!-\!(d-\theta-1)TS\!-\!(d-\theta+z-3)\Phi Q\big{)}\frac{dL}{L}+\left(z-1-\frac{\theta}{d-1}\right)\Phi Q\frac{dr_{0}}{r_{0}}\,,$ (48)
using (39)-(41) to replace $\Theta$, $U$ and $U_{2}$ with the thermodynamic variables $M,T,S,\Phi$ and $Q$. We checked this matches with the extended boundary first law (20) if the pressure satisfies the HVL equation of state and the chemical potential obeys the holographic Euler equation. The match follows from the same holographic dictionary for the thermodynamic variables as used in the main text, (21) and (23)-(26). Hence, this strengthens the holographic dictionary derived in the main text. We do note, however, that the boundary theory seems to organize the thermodynamic variables in a more efficient way, since the boundary energy depends on four other variables, $\tilde{E}=\tilde{E}(\tilde{S},\tilde{Q},V,C)$, whereas the bulk mass seems to depend on five variables $M=M(S,Q,P,r_{F},r_{0}^{-\gamma})$ or $M=M(r_{h},q,L,r_{F},r_{0}^{-\gamma})$. Thus, there seems to be a certain redundancy in the bulk thermodynamic description.
Appendix B Smarr formula and Euler equation for planar geometries
For HVL black branes (with planar horizon topology, $k=0$) the mass satisfies one additional, independent scaling relation
$M(\alpha^{d-\theta-1}S,\alpha^{d-\theta+z-2}Q,P,r_{F},r_{0}^{-\gamma})=\alpha^{d-\theta+z-1}M(S,Q,P,r_{F},r_{0}^{-\gamma})\,,$ (49)
from which the following Smarr relation can be derived [35, 33]
$(d-\theta+z-1)M=(d-\theta-1)TS+(d-\theta+z-2)\Phi Q\,.$ (50)
Together with (41) this implies
$(z+1)M=-2\Theta P+\Phi Q\,.$ (51)
This Smarr-like formula can, in turn, be derived from the scaling relation
$M(S,\alpha^{-1}Q,\alpha^{-2}P,r_{F},r_{0}^{-\gamma})=\alpha^{-(z+1)}M(S,Q,P,r_{F},r_{0}^{-\gamma})\,.$ (52)
This scaling relation follows from combining (38) and (49), so there is only one new scaling relation in addition to the three scaling relations in the previous appendix. The scaling relation (49) arises from rescaling $r_{h}\to\alpha r_{h}$ and $q\to\alpha^{d-\theta+z-2}q$, while (52) follows from sending $L\to\alpha L$.
In the holographic field theory, the thermal state that is dual to the black brane lives on the plane. Here as well there are extra scaling relations for the thermodynamic quantities, namely
$\tilde{E}\propto\tilde{T}^{\frac{d-\theta+z-1}{z}}\,,\quad\tilde{S}\propto\tilde{T}^{\frac{d-\theta-1}{z}}\,,\quad\tilde{Q}\propto\tilde{T}^{\frac{d-\theta+z-2}{z}}\,.$ (53)
This implies the additional scaling relation
$\tilde{E}(\alpha^{d-\theta-1}\tilde{S},\alpha^{d-\theta+z-2}\tilde{Q},V,C)=\alpha^{d-\theta+z-1}\tilde{E}(\tilde{S},\tilde{Q},V,C)\,,$ (54)
which yields
$(d-\theta+z-1)\tilde{E}=(d-\theta-1)\tilde{T}\tilde{S}+(d-\theta+z-2)\tilde{\Phi}\tilde{Q}\,.$ (55)
This is obviously dual to the additional Smarr relation (50) in the bulk, upon invoking the holographic dictionary developed in the main text.
Now, if we combine formula (55) with the holographic Euler equation (2) (which is still valid on the plane) and the Lifshitz, hyperscaling violating equation of state (28), we find the relation
$\mu C=\frac{z-1}{d-\theta-1}\tilde{\Phi}\tilde{Q}-\frac{z}{z-\theta/(d-1)}pV\,.$ (56)
In other words, the boundary grand canonical free energy $\tilde{F}=\tilde{E}-\tilde{T}\tilde{S}-\tilde{\Phi}\tilde{Q}$ (that equals $\mu C$ due to the holographic Euler equation) is not proportional to $pV$, but contains an additional $\tilde{\Phi}\tilde{Q}$ term. If we insert this relation back into the holographic Euler equation, then we find an unusual thermodynamic relation
$\tilde{E}=\tilde{T}\tilde{S}+\frac{d-\theta+z-2}{d-\theta-1}\tilde{\Phi}\tilde{Q}-\frac{z}{z-\theta/(d-1)}pV\,.$ (57)
For $z=1,\theta=0$ this reduces to the standard thermodynamic Euler equation $\tilde{E}=\tilde{T}\tilde{S}+\tilde{\Phi}\tilde{Q}-pV$, that is valid for a thermal CFT on a plane. Also for $\theta=0$ and $\tilde{Q}=0$, equation (57) simplifies to the standard Euler equation $\tilde{E}=\tilde{T}\tilde{S}-pV.$ However, for arbitrary $z$, $\theta$ the generalized Euler equation that involves a $pV$ term seems to be (57) (in contrast to what was claimed in Appendix C of [11] for the case of $\theta=0$ and $\tilde{Q}\neq 0$).
Appendix C Euler equation from a generalized Cardy formula
In $(1+1)$ dimensions the large-$C$ Euler equation can be obtained purely from the field theory side without invoking holography. For two-dimensional CFTs it was shown in [11] that the Euler equation follows from the Cardy formula for the microcanonical entropy [45]. Here we extend that proof to Lifshitz and hyperscaling violating theories using a generalized Cardy formula derived in [46, 47]. This generalized Cardy formula is based on a generalized notion of modular invariance, which allows one to project the partition function at high temperatures to its vacuum state. For Lifshitz theories this invariance has been criticised in [47], since Lifshitz symmetries do not allow one to swap thermal and spatial cycles because it does not contain boosts (for hyperscaling theories this is not an issue since they are Lorentz boost invariant). However, below we proceed with analyzing the implications of the generalized modular invariance and Cardy formula as if it exists.
For arbitrary $(z,\theta)$ the generalized Cardy formula for vanishing angular momentum reads [46, 47, 48]
$\tilde{S}=\frac{1+z-\theta}{1-\theta}2\pi\tilde{E}R^{\frac{z-\theta}{1-\theta}}\left(\frac{-\tilde{E}_{\text{vac}}(1-\theta)}{z\tilde{E}}\right)^{\frac{z}{1-\theta+z}}\,.$ (58)
We have reinstated the radius $R$ of the spatial circle on which the field theory lives, such that the energy is dimensionful and the product $\tilde{E}R^{\frac{z-\theta}{1-\theta}}$ is scale invariant. (For $\theta=0$ the scaling $S\propto R^{z}$ is different from (9) in [46], where $S\propto R$, but our formula seems to have the correct dimensions.) Note the Cardy formula is expressed in terms of the (negative) vacuum energy $\tilde{E}_{\text{vac}}$ instead of the central charge as usual. We can reintroduce a (positive and dimensionless) central charge by defining it in terms of the vacuum Casimir energy
$C:=-\tilde{E}_{\text{vac}}R^{\frac{z-\theta}{1-\theta}}\,.$ (59)
As an example, for two-dimensional CFTs we normalize our central charge as $C=c/12$, where $c$ is the standard central charge appearing in the Casimir energy. For $z=1$ and $\theta=0$ equation (58) reduces to the Cardy formula $\tilde{S}=4\pi R\sqrt{-\tilde{E}\tilde{E}_{\text{vac}}}$ or $\tilde{S}=4\pi\sqrt{\frac{c}{6}L_{0}}$, where we used $\tilde{E}R=2L_{0}$ and $\tilde{E}_{\text{vac}}R=-c/12$ to obtain the latter, more familiar Cardy expression (for $L_{0}=\bar{L}_{0}$).
Since the Cardy formula expresses the entropy as a function $\tilde{S}=\tilde{S}(\tilde{E},V,C),$ with central charge $C$ defined in (59) and the “volume” of the circle given by $V=2\pi R$, we can explicitly compute the temperature, pressure and chemical potential associated with the central charge using the definitions that follow from the first law (20)
$\tilde{T}$ $:=\!\left(\frac{\partial\tilde{S}}{\partial\tilde{E}}\right)_{V,C}^{-1}\!=\!\frac{1}{2\pi R^{\frac{z-\theta}{1-\theta}}}\!\left(\frac{-\tilde{E}_{\text{vac}}(1-\theta)}{z\tilde{E}}\right)^{-\frac{z}{1+z-\theta}},$ (60)
$p$ $:=T\left(\frac{\partial\tilde{S}}{\partial V}\right)_{\tilde{E},C}=\frac{E}{V}\frac{z-\theta}{1-\theta}\,,$ (61)
$\mu$ $:=-T\left(\frac{\partial\tilde{S}}{\partial C}\right)_{\tilde{E},V}=-\frac{\tilde{E}}{C}\frac{z}{1-\theta}\,.$ (62)
In terms of the temperature and central charge the generalized Cardy formula (58) can be written as
$\tilde{S}=\frac{1+z-\theta}{z}(2\pi)^{\frac{1+z-\theta}{z}}C\left(R^{\frac{z-\theta}{1-\theta}}\tilde{T}\right)^{\frac{1-\theta}{z}}\,.$ (63)
Multiplying these thermodynamic variables with their respective conjugate quantities yields
$\tilde{T}\tilde{S}=\tilde{E}\frac{1+z-\theta}{1-\theta}\,,\qquad\,\,pV=\tilde{E}\frac{z-\theta}{1-\theta}\,,\qquad\,\,\mu C=-\tilde{E}\frac{z}{1-\theta}\,.$ (64)
The second relation is the equation of state for a two-dimensional hyperscaling violating theory with Lifshitz symmetry. Moreover, the first and third relation together form the large-$C$ Euler equation (in the absence of charge)
$\tilde{E}=\tilde{T}\tilde{S}+\mu C\,.$ (65)
Hence the Cardy formula knows about the large-$C$ Euler equation! Note the Euler equation does not depend explicitly on $z$ and $\theta$. However, if we replace the $\mu C$ term by a $pV$ term then the formula does depend explicitly on $z$ and $\theta$ (consistent with (57))
$\tilde{E}=\tilde{T}\tilde{S}-\frac{z}{z-\theta}pV\,.$ (66)
Therefore, the large-$C$ Euler equation (65) seems to be a universal thermodynamic relation in large-$C$ field theories, in contrast to (66), which depends on the symmetries of the theory and the number of dimensions.

### v2

@@
 An important feature of this correspondence is that the thermodynamics of stationary AdS black holes [6] is dual to the thermodynamics of high-energy equilibrium states in a strongly coupled, large-$N$ conformal gauge theory [7].
-For instance, the black hole entropy $S=A/4G$ [8, 9] is identified with the thermodynamic entropy $\tilde{S}$ of the CFT equilibrium state, and the internal energy $\tilde{E}$, temperature $\tilde{T}$, electric potential $\tilde{\Phi}$ and charge $\tilde{Q}$ of the CFT state are proportional to, respectively, the mass $M$, temperature $T$, electric potential $\Phi$ and charge $Q$ of the black hole [10, 11, 12, 13]
+For instance, the black hole entropy $S=A/4G$ [8, 9] is identified with the thermodynamic entropy $\tilde{S}$ of the CFT equilibrium state, and the internal energy $\tilde{E}$, temperature $\tilde{T}$, electric potential $\tilde{\Phi}$, and charge $\tilde{Q}$ of the CFT state are proportional to, respectively, the mass $M$, temperature $T$, electric potential $\Phi$ and charge $Q$ of the black hole [10, 11, 12, 13]
 $\tilde{E}=\frac{ML}{R}\,,\,\,\tilde{T}=\frac{TL}{R}\,,\,\,\tilde{S}=\frac{A}{4G}\,,\,\,\tilde{\Phi}=\frac{\Phi}{R}\,,\,\,\tilde{Q}=QL\,$, (1)
@@
 The proportionality factors in the holographic dictionary for $\tilde{E},\tilde{T}$ and $\tilde{\Phi}$ appear due to a specific choice of CFT metric, for which the CFT time is $R/L$ times the static AdS time and the curvature radius is $R$ [4, 5, 14].
-The general principles behind holography are expected to apply to non-asymptotically AdS spacetimes as well.
+The general principles behind holography are expected to apply to nonasymptotically AdS spacetimes as well.
 Since the isometries of the bulk spacetime should match with the global symmetries of the dual boundary theory, field theories that are not conformally (or scale) invariant are not dual to AdS spacetime, but possibly to a different bulk geometry.
@@
 Together with time and space translations and space rotations, this scaling symmetry forms the Lifshitz symmetries.
-For $z\neq 1$, Lifshitz invariant theories describe non-relativistic quantum critical systems.
+For $z\neq 1$, Lifshitz invariant theories describe nonrelativistic quantum critical systems.
 The anisotropic scaling symmetry can be geometrically realized in the bulk by the so-called Lifshitz spacetimes, which are a generalization of AdS spacetime [15, 16, 17, 18] (see [19] for a review).
@@
 We find a generalization of the standard central charge dictionary $C\propto L^{d-1}/G$ for Einstein gravity, and obtain a dual relation between a generalized Smarr formula, relating the thermodynamic quantities in the bulk, and the holographic Euler relation on the boundary [11]
-$\tilde{E}=\tilde{T}\tilde{S}+\tilde{\Phi}\tilde{Q}+\mu C\,,$ (2)
-where the central charge $C$ is defined as the dimensionless proportionality factor of the grand canonical free energy, and $\mu$ is its conjugate chemical potential (we suppress the tildes on $\mu,C$ since they do not have an equivalent definition in the bulk).
-This Euler relation is valid for any large-$C$ field theory at finite temperature in the deconfined phase and it is the same for theories with different global symmetries (e.g., conformal and Lifshitz).
-Moreover, we show that an extended version of the first law for these black holes is dual to the thermodynamic first law on the boundary
+$\tilde{E}=\tilde{T}\tilde{S}+\tilde{\Phi}\tilde{Q}+\mu C\,.$ (2)
+The central charge $C$ is defined as the dimensionless proportionality factor of the energy, entropy and electric charge in the canonical ensemble, i.e., $\tilde{\chi}=Cf_{\tilde{\chi}}(\tilde{T},\tilde{Q},V)$ for $\tilde{\chi}\in\{\tilde{E},\tilde{S},\tilde{Q}\}$.
+Furthermore, the chemical potential is defined as the conjugate quantity to $C$ in the first law, as $\mu\equiv(\partial\tilde{E}/\partial C)_{\tilde{S},\tilde{Q},V}.$ We suppress the tildes on $\mu,C$ since they do not have an equivalent definition in the bulk.
+The Euler relation (2) follows from the nonextensive scaling relation $\tilde{E}(\alpha\tilde{S},V,\alpha\tilde{Q},\alpha C)=\alpha\tilde{E}(\tilde{S},V,\tilde{Q},C)$, where $V$ is the spatial volume.
+This scaling relation only holds in the large-$C$ limit, since there are $1/C$ corrections at finite central charge.
+More specifically, the Euler relation is valid for any large-$C$ field theory at finite temperature in the deconfined phase.
+Moreover, it is the same for theories with different global symmetries, e.g., conformal and Lifshitz symmetries.
+Furthermore, we show that an extended version of the first law for these black holes is dual to the thermodynamic first law on the boundary
 $d\tilde{E}=\tilde{T}d\tilde{S}+\tilde{\Phi}d\tilde{Q}-pdV+\mu dC\,$, (3)
@@
 A line element of the general form [^1] Note that we are working with $d+1$ bulk spacetime dimensions, whereas [33] has $d+2$ bulk dimensions.
-Moreover, we redefined $V_{0}$ in [33] as $-2\Lambda_{0}$, and $\phi_{0}$ as $-\gamma\log r_{0}$, see (11).
+Moreover, we redefined $V_{0}$ in [33] as $2\Lambda_{0}$, and $\phi_{0}$ as $\gamma\log r_{0}$, see (11).
 $ds^{2}$ $=\chi\bigg{[}-\left(\frac{r}{L}\right)^{2z}f(r)dt^{2}+\frac{L^{2}}{f(r)r^{2}}dr^{2}+r^{2}d\Omega_{k,d-1}^{2}\bigg{]}\,$,
 $\chi(r)$ $:=\left(\frac{r}{r_{F}}\right)^{-\frac{2\theta}{d-1}}\,$, (4)
-in $(d+1)$-dimensions, with $f(r)\rightarrow 1$ as $r\rightarrow\infty$, has been suggested to be dual to states in Lifshitz invariant field theories with hyperscaling violation.
+in $(d+1)$ dimensions, with $f(r)\rightarrow 1$ as $r\rightarrow\infty$, has been suggested to be dual to states in Lifshitz invariant field theories with hyperscaling violation.
 For $z=1$ and $\theta=0$ we recover asymptotically AdS spacetime.
-Here, $k$ parameterizes the topology of the $(t,r)=$ constant surfaces: $k=\{-1,0,1\}$ for hyperbolic, planar and spherical topology respectively.
+Here, $k$ parametrizes the topology of the $(t,r)=$ constant surfaces: $k=\{-1,0,1\}$ for hyperbolic, planar and spherical topology respectively.
 When $f(r)=1$ under the rescalings
@@
 Leaving aside the scaling of the bulk radius $r$, this rescaling of the boundary metric is characteristic of hyperscaling violating field theories.
-Hence, we refer to (4) as asymptotically Lifshitz and hyperscaling violating geometries (although, strictly speaking, this is only true for $k=0$, since geometries with $k=\pm 1$ are not invariant under spatial translations.) As is standard for hyperscaling violating theories, the dual black hole geometry is considered to be valid up to a certain large radius $r_{F}$, which corresponds to a UV cut-off scale on the boundary, since the dual field theory could have a different UV completion [28, 31, 34].
-[…]
+Hence, we refer to (4) as asymptotically Lifshitz and hyperscaling violating geometries (although, strictly speaking, this is only true for $k=0$, since geometries with $k=\pm 1$ are not invariant under spatial translations).
+As is standard for hyperscaling violating theories, the dual black hole geometry is considered to be valid up to a certain large radius $r_{F}$, which corresponds to a UV cutoff scale on the boundary, since the dual field theory could have a different UV completion [28, 31, 34].
+Static, charged black holes with line element (4) were found as analytic solutions to a generalized Einstein-Maxwell-Dilaton theory [33], consisting of Einstein gravity minimally coupled to a real scalar field $\phi$, called the “dilaton,” and three Abelian gauge fields $\textsf{A},\textsf{B}$ and C.
+The action of the generalized EMD theory is given by
 $I$ $=\frac{1}{16\pi G}\int d^{d+1}x\sqrt{-g}\Big{[}R-\frac{1}{2}(\nabla\phi)^{2}+\mathcal{V}(\phi)$
@@
 $\mathcal{V}=-2\Lambda_{0}e^{\lambda_{0}\phi}\,,\ X=X_{0}e^{\lambda_{1}\phi}\,,\ Y=Y_{0}e^{\lambda_{2}\phi}\,,\ Z=Z_{0}e^{\lambda_{3}\phi}\,$,
-where the $\lambda_{i}$ are given in terms of ($d,\theta,z)$; namely:
+where the $\lambda_{i}$ are given in terms of ($d,\theta,z)$; namely,
 $\lambda_{0}$ $=\frac{2\theta}{\gamma(d-1)}\,,\quad\lambda_{1}=-\frac{2(d-\theta-1+\theta/(d-1))}{\gamma}\,$,
@@
 $\gamma=$ $\sqrt{2(d-\theta-1)\left(z-1-\frac{\theta}{d-1}\right)}\,$, (8)
-allowing for a non-zero hyperscaling violating parameter $\theta$.
-The gauge field A is responsible for the anisotropic scaling symmetry, B supports non-trivial (spherical or hyperbolic) horizon topology, and C allows for non-zero electric charge.
+allowing for a nonzero hyperscaling violating parameter $\theta$.
+The gauge field A is responsible for the anisotropic scaling symmetry, B supports nontrivial (spherical or hyperbolic) horizon topology, and C allows for nonzero electric charge.
 The analytic solution for these black holes is given by (4) with [33]
@@
 $\phi$ $=\gamma\log(r/r_{0})\;,\quad$ (11)
-generalizing previously discovered solutions for charged spherical black holes with Lifshitz asymptotics [27] (to non-zero $\theta$) and charged black branes with arbitrary Lifshitz and hyperscaling violation parameters [29] (to non-trivial topology).
+generalizing previously discovered solutions for charged spherical black holes with Lifshitz asymptotics [27] (to nonzero $\theta$) and charged black branes with arbitrary Lifshitz and hyperscaling violation parameters [29] (to nontrivial topology).
 The respective mass and charge parameters are $m$ and $q$, and
@@
 The Hawking temperature $T$, Bekenstein-Hawking entropy $S$, ADM mass $M$, electric charge $Q$, and electric potential $\Phi$ are [^2] The ADM mass was obtained in [33] through the background subtraction method, so that the solution with $m=0$ has zero total gravitational energy.
-For hyperbolic black holes, however, this solution does not correspond to the ground state of the grand canonical ensemble, which is given by the extremal solution with zero charge and in our conventions hence has negative energy (whereas in [33] the ADM mass for $k=-1$ was redefined such that the extremal solution has zero energy).
+For hyperbolic black holes, however, this solution does not correspond to the ground state of the grand canonical ensemble, which is given by the extremal solution with zero charge and in our conventions hence has negative energy (whereas in [33] the ADM mass for k=1k=-1 was redefined such that the extremal solution has zero energy).
 $T$ $=\frac{f^{\prime}(r_{h})}{4\pi}\Bigl{(}\frac{r_{h}}{L}\Bigr{)}^{z+1}\!\!=\frac{r_{h}^{z}}{4\pi L^{z+1}}\left[d-\theta+z-1+\frac{k(d-2)^{2}}{d-\theta+z-3}\frac{L^{2}}{r_{h}^{2}}-\frac{(d-\theta+z-3)q^{2}}{r_{h}^{2(d-\theta+z-2)}}\right]\,,\quad S=\frac{\omega_{k,d-1}}{4G}r_{h}^{d-\theta-1}r_{F}^{\theta}\,$,
@@
 The mass also depends on the length scales $r_{F}$ and $r_{0}$, but we treat them here as fixed parameters, since they do not have a clear thermodynamic interpretation from the bulk and boundary perspective.
-From the mass function (35) the following scaling relation may be inferred:
+From the mass function (35) in the Supplemental Matertial the following scaling relation may be inferred:
 $M(S,Q,P)=\alpha^{2+\theta-d}M(\alpha^{d-\theta-1}S,\alpha^{d-\theta+z-3}Q,\alpha^{-2}P)\,$, (15)
@@
 the conjugate quantity to $P$.
-If $\theta=0$ and $z=1$, $\Theta$ reduces to the expected expression $-r^{d}\omega_{k,d-1}/d$.
-Hence $\Theta$ (see equation (42)) is the HVL generalization of (minus) the “thermodynamic volume”, see e.g.
+If $\theta=0$ and $z=1$, $\Theta$ reduces to the expected expression $r^{d}\omega_{k,d-1}/d$.
+Hence $\Theta$ (see equation (42) in the Supplemental Matertial) is the HVL generalization of (minus) the “thermodynamic volume”, see e.g.
 [39].
 The generalized Smarr relation (16) is new in the literature, as the scaling relation (15) was not properly derived before.
-In fact, in [35, 33, 40] it was assumed that the standard Smarr formula for AdS black holes (given by (16) with $z=1$ and $\theta=0$) also holds HVL black holes.
+In fact, in [35, 33, 40] it was assumed that the standard Smarr formula for AdS black holes (given by (16) with $z=1$ and $\theta=0$) also holds for HVL black holes.
 However this is clearly not the case, since the Smarr relation (16) explicitly involves $z$ and $\theta$.
@@
 $V\propto R^{d-1}\,.$ (21)
-The dictionary for the central charge can be read off from the proportionality constant of the grand canonical free energy $F=M-TS-\Phi Q$.
-Since the mass, entropy and charge all scale with the central charge, $M,S,Q\propto C$, we can directly determine its dictionary entry by looking at, for instance, the Bekenstein-Hawking entropy (Holographic dictionary for Lifshitz and hyperscaling violating black holes).
-The central charge is dimensionless and should only depend on theory parameters, not on the specific details of the state, such as its temperature.
-Therefore, the dictionary for $C$ cannot involve the horizon radius $r_{h}$, which is related to the temperature (Holographic dictionary for Lifshitz and hyperscaling violating black holes).
-Multiplying and dividing the entropy by $L^{d-\theta-1}$, we can write
-$S\propto Cx^{d-\theta-1}\,,$ (22)
-with $x=r_{h}/L$, where we identify the central charge as
+The holographic dictionary follows from requiring (i) that the extended bulk first law (19) exactly matches with the boundary first law (20); (ii) that the chemical potential $\mu$ obeys the holographic Euler equation (2); and (iii) that the pressure $p$ satisfies the equation of state for HVL theories [19, 36]
+$p=\frac{\tilde{E}}{V}\frac{z-\theta/(d-1)}{d-1-\theta}\,.$ (22)
+These three conditions yield the dictionary given below in (24)-(28).
+[^3] We also checked that the matching between the bulk and boundary extended first laws continues to hold if $r_{F}$ and $r_{0}$ are allowed to vary in the bulk (see Supplemental Material).
+We note that this equation of state is a consequence of the scaling relation
+$\tilde{E}(\tilde{S},\alpha^{d-1-\theta}V,\tilde{Q},C)=\alpha^{-z+\frac{\theta}{d-1}}\tilde{E}(\tilde{S},V,\tilde{Q},C)\,$, (23)
+which follows from the anisotropic scaling symmetry and the hyperscaling violating property.
+Differentiating (23) with respect to $\alpha$ and setting $\alpha=1$ yields (22), where the pressure is defined as $p:=-(\partial\tilde{E}/\partial V)_{\tilde{S},\tilde{Q},C}$.
+We now discuss the dictionary for the various thermodynamic parameters in turn.
+First, we find the following dictionary for the central charge
 $C\propto\frac{L^{d-\theta-1}r_{F}^{\theta}}{G}\,.$ (24)
 For $\theta=0$ this dictionary reduces to the well-known result for Einstein gravity.
-It is interesting to note that the dictionary entry (23) depends on $\theta$ but not $z$.
+The central charge is dimensionless and should only depend on theory parameters, not on the specific details of the state, such as its temperature.
+Hence, the dictionary for $C$ does not involve the horizon radius $r_{h}$, which is related to the temperature (Holographic dictionary for Lifshitz and hyperscaling violating black holes).
+Further, it is interesting to note that the right-hand side of (24) depends on $\theta$ but not $z$.
 It is well known that $\theta$ effectively lowers the number of dimensions, which is also apparent from (24).
-Moreover, it is quite natural that (23) involves $r_{F}$, since that corresponds to a UV cut-off in the boundary theory, on which the number of field degrees of freedom depends.
-Likewise, the entropy does not depend on $r_{0}$ and so the entry (23) for $C$ does not, since $S\propto C$.
-The dictionary for the internal energy and temperature follows from symmetry arguments.
-The products $ML$ and $TL$ in the bulk are dimensionless, and so should match with products of boundary quantities that are (from (5)) invariant under the anisotropic scaling symmetry, $\{t,x^{i}\}\to\{\alpha^{z}t,\alpha x^{i}\}$, and the simultaneous hyperscaling transformation $\{t,x^{i}\}\to\alpha^{-\theta/(d-1)}\{t,x^{i}\}$ that undoes the rescaling (6) of the metric at the boundary.
+Moreover, it is quite natural that (24) involves $r_{F}$, since that corresponds to a UV cutoff in the boundary theory, on which the number of field degrees of freedom depends.
+We note the dictionary (24) is consistent with the definition of $C$ as the proportionality constant of the energy, entropy and electric charge, i.e., $\tilde{E},S,\tilde{Q}\propto C$.
+For instance, the Bekenstein-Hawking entropy (Holographic dictionary for Lifshitz and hyperscaling violating black holes) is proportional to the central charge,
+$S\propto Cx^{d-\theta-1}\,$, (25)
+where we defined $x\equiv r_{h}/L$. Since $S\propto C$ and the entropy does not depend on $r_{0}$, the entry (24) for $C$ likewise does not.
+Second, the dictionary for the internal energy and temperature follows partly from symmetry arguments.
+We can make the internal energy and temperature dimensionless by simply multiplying them with the bulk curvature radius $L$.
+On the field theory side, this should match with products of boundary quantities that are (from (5)) invariant under the anisotropic scaling symmetry, $\{t,x^{i}\}\to\{\alpha^{z}t,\alpha x^{i}\}$, and the simultaneous hyperscaling transformation $\{t,x^{i}\}\to\alpha^{-\theta/(d-1)}\{t,x^{i}\}$ that undoes the rescaling (6) of the metric at the boundary.
 The invariant combinations are
 $\tilde{E}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}=ML\,,\quad\tilde{T}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}=TL\,.$ (26)
-Note that the scaling dimension of energy and temperature is $[\tilde{E}]=[\tilde{T}]=-[t]=z-\theta/(d-1)$, and the scaling dimension of the boundary curvature scale and spatial coordinate is $[R]=[x^{i}]=-1+\theta/(d-1)$ (see Appendix D in [36]), so that the LHS of (24) is indeed dimensionless.
-The dictionary entries for the electric potential and charge are more complicated, but also essentially follow from dimensional analysis.
+Note that the scaling dimension of energy and temperature is $\tilde{E}]=[\tilde{T}]=-[t]=z-\theta/(d-1)$, and the scaling dimension of the boundary curvature scale and spatial coordinate is $R]=[x^{i}]=-1+\theta/(d-1)$ (see Appendix D in [36]), so that the lhs of (26) is indeed dimensionless.
+Next, the dictionary entries for the electric potential and charge are more complicated, but also follow for a large part from dimensional analysis.
 We take the boundary electric charge $\tilde{Q}$ to be dimensionless, as for the entropy and central charge in (20); the boundary potential $\tilde{\Phi}$ then has the same dimensions as $\tilde{E}$ and $\tilde{T}$, so that $\tilde{\Phi}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}$ is scale invariant.
-The bulk charge scales as $Q\propto qL^{-1}r_{F}^{\theta-\theta/(d-1)}r_{0}^{1-z+\theta/(d-1)}/G$ and the bulk potential as $\Phi\propto qr_{F}^{\theta/(d-1)}r_{0}^{z-1-\theta/(d-1)}/(L^{z}r_{h}^{d-\theta+z-3})$, where the charge parameter and Newton’s constant have mass dimension $[q]=-d+\theta-z+2$ and $[G]=-d+1$, respectively.
+The bulk charge scales as $Q\propto qL^{-1}r_{F}^{\theta-\theta/(d-1)}r_{0}^{1-z+\theta/(d-1)}/G$ and the bulk potential as $\Phi\propto qr_{F}^{\theta/(d-1)}r_{0}^{z-1-\theta/(d-1)}/(L^{z}r_{h}^{d-\theta+z-3})$, where the charge parameter and Newton’s constant have mass dimension $q]=-d+\theta-z+2$ and [G]=d+1[G]=-d+1, respectively.
 We can thus make the boundary charge $\tilde{Q}$ and the product $\tilde{\Phi}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}$ dimensionless with the following dictionary
@@
 This difference ultimately arises from the fact that in [11] (following [35, 33]) it was assumed that the AdS Smarr formula also holds for Lifshitz black holes, whereas here we find that a more careful study of the scaling relation for the mass shows that the correct Smarr formula is (16).
-Crucially, using the dictionary in (21) and (23)-(26) one can show that the extended bulk first law (19) exactly matches with the boundary first law (20).
-[^3] We also checked that the matching between the bulk and boundary extended first laws continues to hold if $r_{F}$ and $r_{0}$ are allowed to vary in the bulk (see Supplemental Material).
-The matching only works if the pressure satisfies the equation of state for HVL theories [19, 36]
-$p=\frac{\tilde{E}}{V}\frac{z-\theta/(d-1)}{d-1-\theta}\,,$ (28)
-and if the chemical potential $\mu$ obeys the holographic Euler equation (2). The equation of state (28) is a consequence of the scaling relation
-$\tilde{E}(\tilde{S},\alpha^{d-1-\theta}V,\tilde{Q},C)=\alpha^{-z+\frac{\theta}{d-1}}\tilde{E}(\tilde{S},V,\tilde{Q},C)\,,$ (29)
-which follows from the anisotropic scaling symmetry and the hyperscaling violating property.
-Differentiating (29) with respect to $\alpha$ and setting $\alpha=1$ yields (28), where the pressure is defined as $p:=-(\partial\tilde{E}/\partial V)_{\tilde{S},\tilde{Q},C}$.
+We emphasize that the left-hand side of the holographic dictionary (26)-(28) follows from invariance under Lifshitz and hyperscaling transformations.
+However, in order to fix the right-hand side of these equations, we imposed that the boundary and bulk extended thermodynamics should be consistent with each other.
+Specifically, we imposed the three conditions stated above (22).
 On a planar boundary geometry (for $k=0$) there is an additional relation between the thermodynamic quantities, which is dual to an extra Smarr formula in the bulk.
-This allows us to replace the $\mu C$ term in the Euler relation with a $pV$ term, but this comes at the expense of additional $(z,\theta,d)$-dependent prefactors in front of the $pV$ and $\tilde{\Phi}\tilde{Q}$ terms (that are absent for $z=1,\theta=0$), namely:
+This allows us to replace the $\mu C$ term in the Euler relation with a $pV$ term, but this comes at the expense of additional $(z,\theta,d)$-dependent prefactors in front of the $pV$ and $\tilde{\Phi}\tilde{Q}$ terms (that are absent for $z=1,\theta=0$), namely,
 $\tilde{E}=\tilde{T}\tilde{S}+\frac{d-\theta+z-2}{d-\theta-1}\tilde{\Phi}\tilde{Q}-\frac{z}{z-\theta/(d-1)}pV\,.$ (30)
-We refer to the Supplemental Material for more details.
+This implies that the grand canonical free energy $\tilde{F}=\tilde{E}-\tilde{T}\tilde{S}-\tilde{\Phi}\tilde{Q}$ is not equal to pV-pV for generic $z,\theta.$ We refer to the Supplemental Material for more details.
 Finally, we demonstrate that the large-$C$ Euler equation (2) is dual to the generalized Smarr relation (16).
@@
 To obtain this we used the dictionary for $\tilde{Q}$ (28) and $C$ (24). Inserting this back into (Holographic dictionary for Lifshitz and hyperscaling violating black holes) gives
-$2\Theta P$ $=R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}\frac{1}{L}\left((2-z)\tilde{\Phi}\tilde{Q}+(d-\theta-1)\mu C-\tilde{E}\right)\,.$ (34)
+[…]
 This precisely agrees with the generalized Smarr formula (16) if we insert the large-$C$ Euler equation (2) and employ the holographic dictionary (26), (27) and (28).
@@
 For future work, it would be interesting to derive the dictionary for the thermodynamic black hole quantities using holographic renormalization [42, 43, 44] and, in particular, check our formula for the central charge (which can be read off from the Casimir energy for even $d$ and $k\neq 0$).
-Moreover, we would like to understand whether there exists a thermodynamic interpretation of the length parameters $r_{F}$ and $r_{0}$.
+Moreover, we would like to understand whether there exists a thermodynamic interpretation of the length parameters $r_{F}$ and $r_{0}$, perhaps relating these to “Lifshitz charges”, as proposed in a simplified setting in [36].
 Finally, it is likely that our new setup will lead to a discovery of so far unseen phase transitions and critical phenomena arising for Lifshitz and hyperscaling violating theories.
@@
 D.K.
-is grateful for support from GAČR 23-07457S grant of the Czech Science Foundation and the Charles University Research Center Grant No.
+is grateful for support from Grant No.
+GAČR 23-07457S of the Czech Science Foundation and the Charles University Research Center Grant No.
 UNCE24/SCI/016.
 M.R.V.
-acknowledges support from the SNF Postdoc Mobility grant P500PT-206877 and the Spinoza Grant of the Dutch Science Organisation (NWO).
+acknowledges support from the Swiss National Science Foundation Postdoc Mobility Grant No.
+P500PT-206877 and the Spinoza Grant of the Dutch Science Organisation (NWO).
 R.B.M.
@@
 $M$ $=(d-1-\theta)\frac{r_{F}^{\theta}\omega_{k,d-1}}{16\pi G}\bigg{(}\frac{16\pi GP}{(d-2+z-\theta)(d-1+z-\theta)}\bigg{)}^{\frac{z+1}{2}}\bigg{(}\frac{4GS}{\omega_{k,d-1}r_{F}^{\theta}}\bigg{)}^{\frac{d-1+z-\theta}{d-1-\theta}}\bigg{(}\frac{r_{0}}{r_{F}}\bigg{)}^{-\frac{(z+1)\theta}{d-1}}\times$
-×$\quad\times\bigg{(}1+\frac{(d-2)^{2}(d-\theta+z-2)(d-\theta+z-1)k}{16\pi GP(d-\theta+z-3)^{2}}\left(\frac{r_{0}}{r_{F}}\right)^{\frac{2\theta}{d-1}}\left(\frac{4GS}{\omega_{k,d-1}r_{F}^{\theta}}\right)^{-\frac{2}{d-\theta-1}}$
+$\quad\times\bigg{(}1+\frac{(d-2)^{2}(d-\theta+z-2)(d-\theta+z-1)k}{16\pi GP(d-\theta+z-3)^{2}}\left(\frac{r_{0}}{r_{F}}\right)^{\frac{2\theta}{d-1}}\left(\frac{4GS}{\omega_{k,d-1}r_{F}^{\theta}}\right)^{-\frac{2}{d-\theta-1}}$
 $\qquad+\frac{8\pi GQ^{2}r_{0}^{2(z-1)}(d-\theta+z-2)(d-\theta+z-1)}{PZ_{0}r_{F}^{2\theta}(d-\theta-1)(d-\theta+z-3)\omega_{k,d-1}^{2}}\left(\frac{4GS}{\omega_{k,d-1}r_{F}^{\theta}}\right)^{-\frac{2(d-\theta+z-2)}{d-\theta-1}}\bigg{)}\,.$ (35)
@@
 Using (35) it is straightforward to directly verify the following three independent scaling relations:
-$M(S,Q,P,r_{F},r_{0}^{-\gamma})\,$ $=\alpha^{-\theta}M(\alpha^{\theta}S,\alpha^{\theta-\theta/(d-1)}Q,\alpha^{-2\theta/(d-1)}P,\alpha r_{F},r_{0}^{-\gamma})\quad$ (38)
+$M(S,Q,P,r_{F},r_{0}^{-\gamma})\,$ $=\alpha^{-\theta}M(\alpha^{\theta}S,\alpha^{\theta-\theta/(d-1)}Q,\alpha^{-2\theta/(d-1)}P,\alpha r_{F},r_{0}^{-\gamma})\quad\$(38)
 $=M(S,\alpha^{\lambda_{3}/2}Q,\alpha^{-\lambda_{0}}P,r_{F},\alpha r_{0}^{-\gamma})$
@@
 $\Theta:=-\frac{\partial M}{\partial P}\Big{|}_{S,Q,r_{F},r_{0}}$ $=-\frac{L^{1-z}r_{F}^{\frac{(d+1)\theta}{d-1}}r_{0}^{-\gamma\lambda_{0}}r_{h}^{d-\theta+z-3}(d-\theta-1)\omega_{k,d-1}}{2(d-\theta+z-3)^{2}(d-\theta+z-2)(d-\theta+z-1)}\times$ (42)
-×$\quad\times\bigg{(}r_{h}^{-2(d+z-1)}(d-\theta+z-3)^{2}\big{(}(z+1)r_{h}^{2(d+z)}+q^{2}(z-1)r_{h}^{2\theta+4}\big{)}+(d-2)^{2}kL^{2}(z-1)\bigg{)}\,,$
+$\quad\times\bigg{(}r_{h}^{-2(d+z-1)}(d-\theta+z-3)^{2}\big{(}(z+1)r_{h}^{2(d+z)}+q^{2}(z-1)r_{h}^{2\theta+4}\big{)}+(d-2)^{2}kL^{2}(z-1)\bigg{)}\,$,
 $U:=\frac{\partial M}{\partial r_{F}}\Big{|}_{S,Q,P,r_{0}}$ $=-\frac{\theta\omega_{k,d-1}L^{-z-1}r_{F}^{\theta-1}r_{h}^{-d-\theta-z-3}}{16\pi G(d-1)(d-\theta+z-3)^{2}}\bigg{[}r_{h}^{2}(d-\theta+z-3)^{2}\bigg{(}q^{2}r_{h}^{2\theta+4}((1-2z)d+\theta+(\theta+2)z-1)$
@@
 $\quad+r_{h}^{2}(d-\theta+z-3)^{2}\left(\lambda_{0}(z+1)r_{h}^{2(d+z)}+q^{2}r_{h}^{2\theta+4}(\lambda_{0}(z-1)-2\lambda_{3})\right)\bigg{]}\,.$ (44)
-The quantity $U$ vanishes if $\theta=0$, and $U_{2}$ is zero if both $\theta=0$ and $z=1$, in which case $\Theta$ reduces to $-r^{d}\omega_{k,d-1}/d$.
+The quantity $U$ vanishes if $\theta=0$, and $U_{2}$ is zero if both $\theta=0$ and $z=1$, in which case $\Theta$ reduces to $r^{d}\omega_{k,d-1}/d$.
 It is possible to get rid of the explicit $z$ and $\theta$ dependence in the Smarr relation, but this comes at the expense of introducing two additional terms, proportional to $r_{F}$ and $r_{0}^{-\gamma}$, yielding the generalized Smarr formula
@@
 $dM=TdS+\Phi dQ-\Theta dP+Udr_{F}+U_{2}dr_{0}^{-\gamma}\,.$ (47)
-Since the pressure depends via $\Lambda_{0}$ on $L$, $r_{F}$ and $r_{0}$, due to (12), the extended first law can also be written in terms of variations of these length scales
+Since the pressure depends via $\Lambda_{0}$ on $L$, $r_{F}$ and $r_{0}$, due to (12) in the main text, the extended first law can also be written in terms of variations of these length scales
 $dM=TdS+\Phi dQ+\big{(}\theta M-\theta TS-\frac{(d-2)\theta}{d-1}\Phi Q\big{)}\frac{dr_{F}}{r_{F}}$
@@
 For $z=1$ and $\theta=0$ equation (58) reduces to the Cardy formula $\tilde{S}=4\pi R\sqrt{-\tilde{E}\tilde{E}_{\text{vac}}}$ or $\tilde{S}=4\pi\sqrt{\frac{c}{6}L_{0}}$, where we used $\tilde{E}R=2L_{0}$ and $\tilde{E}_{\text{vac}}R=-c/12$ to obtain the latter, more familiar Cardy expression (for $L_{0}=\bar{L}_{0}$).
-Since the Cardy formula expresses the entropy as a function $\tilde{S}=\tilde{S}(\tilde{E},V,C),$ with central charge $C$ defined in (59) and the “volume” of the circle given by $V=2\pi R$, we can explicitly compute the temperature, pressure and chemical potential associated with the central charge using the definitions that follow from the first law (20)
+Since the Cardy formula expresses the entropy as a function $\tilde{S}=\tilde{S}(\tilde{E},V,C),$ with central charge $C$ defined in (59) and the “volume” of the circle given by $V=2\pi R$, we can explicitly compute the temperature, pressure and chemical potential associated with the central charge using the definitions that follow from the first law
 $\tilde{T}$ $:=\!\left(\frac{\partial\tilde{S}}{\partial\tilde{E}}\right)_{V,C}^{-1}\!=\!\frac{1}{2\pi R^{\frac{z-\theta}{1-\theta}}}\!\left(\frac{-\tilde{E}_{\text{vac}}(1-\theta)}{z\tilde{E}}\right)^{-\frac{z}{1+z-\theta}},$ (60)